Suppose the market for cotton is competitive. A typical cotton farmer has a total cost function...

  1. Suppose the market for cotton is competitive. A typical cotton farmer has a total cost

    function of:

    C = 100 + 15q – 6q2 + q3. The prevailing market price is $15.

    1. Find the profit-maximizing output level of this farmer. Calculate the corresponding profit at this output level. Show your steps. (7 marks)

    2. Suppose all the fixed cost is unavoidable. Explain whether this farmer should shut down its production in the short run. (2 marks)
Answer #1

1) Me= da [100+ 1500 - 699 1703] 15 - 12 Q1 + 392 - At man profir P = MC > 15-12q +39? 15 19 4 unite TR = P Q = 15+4=60 TC= 1- TVc:c- 100 = 15 q- 6 atq} AVC = TVC = 15-6 qr q² - u 2 + 15-24 = 7 AUCEP .: Firm should not shut down